# Angular momentum of vacuum bubbles in a first-order phase transition

### v1

@@ 4.1 Evolution of perturbations: $\eta_{0}$ to $\eta_{\rm c}$
The initial conditions are set by the primordial curvature perturbation $\mathcal{R}_{k}(0)$ associated with the comoving wavenumber $k$, when modes are superhorizon. In the case where the background Hubble evolution is determined solely by a component with constant $w$, and with a constant sound speed $c_{s}^{2}=w$, the evolution equations for the perturbations admit exact solutions. For the SM plasma with $w=c_{s}^{2}=1/3$, one can show that the density and velocity perturbations in the UHG are [51]
@@
Observe that the perturbations in the UHG, given by Eqs. (4.13) and (4.15) are nonsingular in the superhorizon regime $z\rightarrow 0$, and this is another advantage of adopting the UHG, as pointed out by, e.g.  [40] and [21]. In fact, as shown in the left panel of Fig. 2, the transfer functions are close to unity at early times, i.e. when the modes are superhorizon. As for $\nabla^{2}\gamma$ in Fourier space, we can obtain it by integrating the ordinary differential equation (ODE),
@@
where we set the initial condition to be such that $-k^{2}\gamma_{k}=0$ at $z=0$.
@@
When the mode enters the horizon, i.e. when $k\sim\mathcal{H}=1/\eta$, the perturbations start to oscillate as they feel the presence of the radiation fluid with a nonzero sound speed. One can similarly write down the evolution equations for $\delta_{\rm D}$ and $\hat{\psi}_{\rm D}$, and find that
@@ 4.2.1 Background evolution
From the critical temperature up to percolation, the usual cosmological evolution of the Universe is modified by first-order phase transition (FOPT) dynamics. The FOPT is triggered mainly by the effective potential given in Eq. (2.2) developing a nonzero vacuum expectation value.
@@
$\rho_{\rm SM}$ $=$ $\rho_{\rm SM,c}a^{-4}\,,\quad\rho_{\rm SM,c}=\frac{\pi^{2}}{30}g_{\rho\rm,SM,c}T_{\rm SM,c}^{4}\,.$ (4.19)
The value of $g_{\rho,\rm SM,c}$ is taken from [52]. The reduced Hubble parameter, which is obtained by normalizing the Hubble parameter relative to its value at the critical temperature, is given by
@@
[…]
The wall velocity is as an important ingredient in tracking the false vacuum fraction. In particular, earlier work by [53, 54, 24] incorporated the evolution of the bubble wall velocity, where reheating effects in the case of deflagrations will slow down the advancing bubble wall. Recent work by, e.g. [55, 56], provided prescriptions to obtain the bubble wall velocity in a model-independent manner. Here we take a simplified approach, following Ref. [57], where
$v_{\rm w}=1-\left(\frac{T}{T_{\rm c}}\right)^{k}\,,\quad k=8\,.$ (4.23)
We have checked that quantities that do not explicitly depend on the wall velocity, such as the percolation temperature and inverse FOPT duration, marginally depend on $k$. However, the spin of FV bubbles has a strong dependence on $v_{\rm w}$. As we remarked in Section 2, the nucleation rate depends mainly on the particle physics model. The exact connection lies in the following expression
@@
where $S_{3}(T)$ is the bounce action associated with the scalar field configuration that satisfies the equation of motion. In the case of quartic effective potentials for a single real scalar field, a semianalytic expression has been formulated in Ref. [58], which we adopt in this work. For our effective potential, we take $S_{3}$ to be
@@
We can now formulate the system of coupled first order ODEs which describe the background evolution in the false vacuum during the FOPT. The ODEs track $x\equiv H_{\rm c}(t-t_{\rm c}),y\equiv T/T_{\rm c},F$, the derivatives of $F$, and $\tilde{\eta}\equiv\mathcal{H}_{\rm c}(\eta-\eta_{\rm c})$. We set $a=1$ at the critical point. This then implies that $H_{\rm c}=\mathcal{H}_{\rm c}$. Taking the scale factor $a$ to be the independent variable, one can show that
@@
[…]
The $g_{n}$s are dimensionless quantities that are related to the derivatives of $\ln F$, and can also be defined recursively via Eq. (4.31). A similar system of ODEs, in particular that converts the integro-differential equation into a system of ODEs, was obtained in Ref. [59] and recently in Ref. [60]. Note that our system of ODEs has one less equation than Ref. [60], if we do not include Eq. (4.33) in the counting, which only tracks the elapsed conformal time from the critical point in units of $H_{\rm c}$. The initial conditions are
@@
for $a=1$. As for the temperature of the SM sector, we take it to be
$T_{\rm SM}\simeq\frac{T_{\rm SM,c}}{a}\simeq\frac{T_{\rm c}}{a}\frac{1}{r_{T\rm,c}}\,,$ (4.36)
where $r_{T\rm,c}$ is the dark to SM temperature ratio at the critical point. Finally, we check that the physical volume in the false vacuum regions is decreasing, at least at the percolation time. This can be expressed as [61, 62]
@@
Figure 3: Evolution of the temperature and FV fraction (top row), and of $c_{s\rm,D}^{2}$ and $w_{\rm D}$ (bottom), for BP-2 (column a) and BP-7 (column b). The vertical black dashed lines indicate the percolation temperature in all panels, while the vertical dot-dashed and dotted lines respectively indicate $\tilde{\eta}_{i}$ and $\tilde{\eta}_{f}$ in the bottom panels.
@@
For illustration purposes, in Fig. 3 and in the following figures, we feature two specific FOPT scenarios: BP-2 and BP-7 from Table 1. For each case, the evolution of the temperature (solid) and false vacuum fraction (dashed) versus the rescaled elapsed conformal time $H_{\rm c}(\eta-\eta_{\rm c})$, are shown in the top row plots. The labeling is based on Table 1, where we list the benchmark FOPT scenarios of interest.
BP-1 BP-2 BP-3 BP-4 BP-5 BP-6 BP-7 BP-8
$A$ $10^{-1}$ $10^{-1}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-3}$ $10^{-4}$
$x_{C}$ $1.0365\times 10^{-2}$ $1.0365\times 10^{-2}$ $2.251\times 10^{-3}$ $2.251\times 10^{-3}$ $2.251\times 10^{-3}$ $1.4989\times 10^{-4}$ $2.1463\times 10^{-5}$ $1.6172\times 10^{-5}$
$\lambda$ 0.1836 0.1836 $1.4043\times 10^{-2}$ $1.4043\times 10^{-2}$ $1.4043\times 10^{-2}$ $8.522\times 10^{-3}$ $3.9556\times 10^{-4}$ $1.836\times 10^{-5}$
$L_{\rm c}/T_{\rm c}^{4}$ $9.0569\times 10^{-2}$ $9.0569\times 10^{-2}$ $5.0496\times 10^{-2}$ $5.0496\times 10^{-2}$ $5.0496\times 10^{-2}$ $4.8887\times 10^{-2}$ $9.0569\times 10^{-2}$ 0.8
$T_{\rm c}$ (MeV) $10^{5}$ $10$ $10^{5}$ $01$ $10$ $10^{3}$ $10^{-2}$
$r_{T\rm,c}$ 0.4 0.4 0.4 $0.1$ 0.4 0.4 0.4 $0.1$
$T_{*}$ (MeV) $8.7132\times 10^{4}$ 8.8996 $8.0128\times 10^{4}$ 8.1911 0.8299 6.6853 $6.5256\times 10^{2}$ $8.788\times 10^{-3}$
$\sqrt{3}H_{*}R_{*}$ $1.1926\times 10^{-3}$ $6.9296\times 10^{-4}$ $1.9432\times 10^{-3}$ $1.4573\times 10^{-3}$ $1.2045\times 10^{-3}$ $2.1579\times 10^{-3}$ $6.5936\times 10^{-4}$ $5.6272\times 10^{-5}$
$M_{*}/M_{\odot}$ $3.2916\times 10^{-19}$ $1.6059\times 10^{-10}$ $1.7103\times 10^{-18}$ $2.4759\times 10^{-13}$ $1.1649\times 10^{-7}$ $8.2059\times 10^{-9}$ $1.2081\times 10^{-15}$ $5.1961\times 10^{-11}$
$\alpha_{*}$ $2.7930\times 10^{-5}$ $2.3249\times 10^{-4}$ $2.5184\times 10^{-5}$ $5.7671\times 10^{-7}$ $2.1285\times 10^{-4}$ $3.1267\times 10^{-4}$ $1.0988\times 10^{-4}$ $4.8694\times 10^{-5}$
$(\beta/H)_{*}$ $1.4602\times 10^{3}$ $2.3087\times 10^{3}$ $1.1229\times 10^{3}$ $1.4446\times 10^{3}$ $1.7025\times 10^{3}$ $1.1835\times 10^{3}$ $3.9481\times 10^{3}$ $5.3696\times 10^{4}$
$v_{\rm w,*}$ 0.6678 0.6065 0.8301 0.7974 0.7751 0.9601 0.9671 0.6443
$s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$ 9.1707 3.0934 3.4005 $2.5764\times 10^{2}$ 0.901 0.3341 $2.7841\times 10^{1}$ $8.6758\times 10^{4}$
$\tilde{\eta}_{*}-\tilde{\eta}_{i}$ $6.8894\times 10^{-3}$ $4.563\times 10^{-3}$ $9.7022\times 10^{-3}$ $7.6396\times 10^{-3}$ $6.531\times 10^{-3}$ $1.1398\times 10^{-2}$ $4.019\times 10^{-3}$ $7.9014\times 10^{3}$
$c_{s\rm,D*}^{2}$ $1.9829\times 10^{-2}$ $1.4258\times 10^{-2}$ $3.2436\times 10^{-2}$ $2.8191\times 10^{-2}$ $2.5533\times 10^{-2}$ $1.6212\times 10^{-2}$ $2.4072\times 10^{-3}$ $4.1637\times 10^{-4}$
$\tilde{\eta}_{*}$ 0.1477 0.1236 0.248 0.2208 0.205 0.4958 0.5324 0.1379
$\mathcal{C}_{*}$ $2.866\times 10^{1}$ $4.5227\times 10^{2}$ 3.9446 $3.7088\times 10^{1}$ $1.3023\times 10^{2}$ 1.8728 $1.2146\times 10^{1}$ $1.0821\times 10^{5}$
$v_{\rm eff,*}/A_{\rm s}$ $4.7884\times 10^{-3}$ $1.1053\times 10^{-2}$ $2.8945\times 10^{-3}$ $6.6561\times 10^{-3}$ $1.0309\times 10^{-2}$ $2.2148\times 10^{-3}$ $1.7235\times 10^{-3}$ $1.3883\times 10^{-2}$
$s_{\rm rms,*}$ $4.3913\times 10^{-2}$ $3.419\times 10^{-2}$ $9.8429\times 10^{-3}$ 1.7149 $9.2887\times 10^{-3}$ $7.4004\times 10^{-4}$ $4.7984\times 10^{-2}$ $1.2045\times 10^{3}$
@@
We mark the moment of percolation with a vertical dashed line. In both cases, we have taken $r_{T\rm,c}=0.4$. For these two cases, the percolation temperatures are such that $T_{*}/T_{\rm c}<0.9$, and the phase transition strength $\alpha_{*}$ is at least $10^{-4}$. Here we define $\alpha_{*}$ as
$\alpha_{*}$$\equiv$ $\frac{1}{\rho_{\rm R,tot}}\left(1-\frac{T}{4}\frac{d}{dT}\right)\Delta V(T)\Bigg{|}_{T=T_{*}},\quad\rho_{\rm R,tot}(T_{*})=\frac{\pi^{2}}{30}T_{*}^{4}\left[g_{\rho}+\frac{g_{\rho,{\rm SM}}(T_{*})}{r_{T\rm,c}^{4}}\right]\,.$ (4.39)
The difference in the conformal time elapsed, from the critical point up to the percolation time, for these two cases can be estimated by first noting that $T\propto 1/a$ and $a\propto\eta$ in the radiation-dominated era. Then
@@
From the above estimate, a lower percolation temperature implies a longer conformal time duration between the critical point and percolation. The percolation temperature, in turn, can be approximated by [63]
[…]
where
$\frac{\beta(t)}{H}=\frac{1}{H}\frac{\dot{\Gamma}}{\Gamma}\simeq T\frac{d}{dT}\left(\frac{S_{3}}{T}\right)$ (4.42)
is the inverse time scale of the FOPT at time $t$. In the last line, we assume that the nucleation rate increases monotonically with time, the scale factor is approximately constant, and $\beta_{*}(t-t_{\rm c})\gg 1$. In the final step, we replaced $\Gamma_{*}$ with $\Gamma(t)$, and $\beta_{*}$ with $\beta(t)$. Then
$1\approx\frac{8\pi}{3}v_{\rm w,*}^{3}\frac{\Gamma_{*}}{\beta_{*}^{4}}\Rightarrow\exp\left[\frac{S_{3}(T_{*})}{T_{*}}\right]\approx\frac{8\pi}{3}\left[\frac{S_{3}(T_{*})}{T_{*}}\right]^{3/2}\frac{v_{\rm w}^{3}}{(\beta_{*}/H_{*})^{4}}\left(\frac{T_{*}}{H_{*}}\right)^{4}\,.$ (4.43)
Typically, the bounce action at percolation for an FOPT with $T_{\rm c}=10\,\mathrm{MeV}$ and $r_{T\rm,c}=0.4$ is $S_{3}/T\sim 150$. In Fig. 4 we plot $S_{3}/T$ as solid curves, where we feature BP-2 (red) and BP-7 (blue). It can be seen that the typical value $(S_{3}/T)\sim 150$ corresponds to a higher percolation temperature for BP-2 compared to BP-7. For reference, we indicated the actual percolation temperatures as vertical dot dashed lines. Note that the actual values of $S_{3}(T_{*})/T_{*}$ for BP-2 and BP-7 are, respectively, 155.78 and 133.86. The difference can be mainly attributed to the fact that these benchmark points have different critical temperatures. In fact, based on the estimate in Eq. (4.43), the respective values of $S_{3}(T_{*})/T_{*}$ for BP-2 and BP-7 are expected to differ by $4\ln\left[\left(1\,\mathrm{GeV}/10\,\mathrm{MeV}\right)\right]\simeq 18$.
Figure 4: $S_{3}/T$ and $\beta/H$, for BP-2 (red) and BP-7 (blue). The vertical dot-dashed lines indicate the percolation temperature for each benchmark.
In order to account for the presence of the FV and TV, the thermodynamic properties of the dark fluid are determined by the average of the FV and TV contributions, weighted by the respective volume fractions, i.e.
@@
We comment that the above prescription, which uses volume fractions as weights to account for the contribution of each phase, is valid over sufficiently large scales where FV and TV features are indistinguishable. Meanwhile, the dark sector equation of state, and the sound speed are then respectively given by
$\boxed{w_{\rm D}=\frac{p_{\rm D}}{\rho_{\rm D}},\quad c_{s\rm,D}^{2}=\frac{dp_{\rm D}/dt}{d\rho_{\rm D}/dt}=\frac{s_{\rm D}-\left(\frac{\dot{F}}{H}\right)\frac{\Delta p}{T}}{T~{}\frac{\partial s_{\rm D}}{\partial T}-\left(\frac{\dot{F}}{H}\right)\frac{\Delta\rho}{T}}=\frac{1}{3}\left[\frac{1-(1-F)\frac{T\Delta s}{4p_{\rm FV}}}{1-\frac{\dot{F}}{H}\frac{\Delta\rho}{4\rho_{\rm FV}}}\right]}\,.$ (4.45)
@@
$w_{\rm FV}$ False vacuum equation of state in the dark sector
$w_{\rm TV}$ True vacuum equation of state in the dark sector
$w_{\rm D}$ Dark sector equation of state, weighted over the false and true vacua
@@
$w$ Average equation of state, weighted over the dark and visible sectors
$c_{s\rm,FV}^{2}$ False vacuum sound speed squared in the dark sector; always equal to $1/3$
$c_{s\rm,TV}^{2}$ False vacuum sound speed squared in the dark sector; not necessarily $1/3$
$c_{s\rm,D}^{2}$ Dark sector sound speed squared, weighted over the false and true vacua

### v2

@@ 4.1 Evolution of perturbations: $\eta_{0}$ to $\eta_{\rm c}$
The initial conditions are set by the primordial curvature perturbation $\mathcal{R}_{k}(0)$ associated with the comoving wavenumber $k$, when modes are superhorizon. In the case where the background Hubble evolution is determined solely by a component with constant $w$, and with a constant sound speed $c_{s}^{2}=w$, the evolution equations for the perturbations admit exact solutions. For the SM plasma with $w=c_{s}^{2}=1/3$, one can show that the density and velocity perturbations in the UHG are [Hwang:1993ai]
@@
Observe that the perturbations in the UHG, given by Eqs. (4.13) and (4.15) are nonsingular in the superhorizon regime $z\rightarrow 0$, and this is another advantage of adopting the UHG, as pointed out by, e.g.[Bardeen:1980kt] and [Schmid:1998mx]. In fact, as shown in Fig. 2, the transfer functions are close to unity at early times, i.e., when the modes are superhorizon. As for $\nabla^{2}\gamma$ in Fourier space, we can obtain it by integrating the ordinary differential equation (ODE),
@@
where we set the initial condition to be such that $k^{2}\gamma_{k}=0$ at $z=0$.
@@
When the mode enters the horizon, i.e., when $k\sim\mathcal{H}=1/\eta$, the perturbations start to oscillate as they feel the presence of the radiation fluid with a nonzero sound speed. One can similarly write down the evolution equations for $\delta_{\rm D}$ and $\hat{\psi}_{\rm D}$, and find that
@@ 4.2.1 Background evolution
From the critical temperature up to percolation, the usual cosmological evolution of the Universe is modified by FOPT dynamics. The FOPT is triggered mainly by the effective potential given in Eq. (2.2) developing a nonzero vacuum expectation value.
@@
$\rho_{\rm SM}$ $=$ $\frac{\pi^{2}}{30}g_{\rho\rm,SM}(T_{\rm SM})T_{\rm SM}^{4}\,,\quad\rho_{\rm SM,c}=\frac{\pi^{2}}{30}g_{\rho\rm,SM,c}T_{\rm SM,c}^{4}\,.$ (4.19)
The value of $g_{\rho,\rm SM}(T_{\rm SM})$ is taken from Ref. [Drees:2015exa]. The reduced Hubble parameter, which is obtained by normalizing the Hubble parameter relative to its value at the critical temperature, is given by
@@
$-\ln F(t)=\frac{4\pi}{3}\int_{t_{\rm c}}^{t}dt^{\prime}~\Gamma(T(t^{\prime}))~a^{3}(t^{\prime})r^{3}(t,t^{\prime})\,,\quad r(t,t^{\prime})\equiv\int_{t^{\prime}}^{t}dt^{\prime\prime}~\frac{v_{\rm w}(t^{\prime\prime})}{a(t^{\prime\prime})}\,.$ (4.22)
The wall velocity is as an important ingredient in tracking the false vacuum fraction. In particular, earlier work [Heckler:1994uu, Megevand:2000da, Leitao:2014pda] incorporated the evolution of the bubble wall velocity, where reheating effects in the case of deflagrations will slow down the advancing bubble wall. Recent work, e.g.[Ai:2021kak, Ai:2023see], provided prescriptions to obtain the bubble wall velocity in a model-independent manner. Here we take a simplified approach, following Ref. [kampfer1988phenomenologicalPaper2], where
$v_{\rm w}=1-\left(\frac{T}{T_{\rm c}}\right)^{n_{\rm w}}\,,\quad n_{\rm w}=8\,.$ (4.23)
We have checked that quantities that do not explicitly depend on the wall velocity, such as the percolation temperature and inverse FOPT duration, marginally depend on $n_{\rm w}$. However, the spin of FV bubbles has a strong dependence on $v_{\rm w}$. As we remarked in Section 2, the nucleation rate depends mainly on the particle physics model. The exact connection lies in the following expression
@@
where $S_{3}(T)$ is the bounce action associated with the scalar field configuration that satisfies the equation of motion. In the case of quartic effective potentials for a single real scalar field, a semianalytic expression has been formulated in Ref. [Adams:1993zs], which we adopt in this work. For our effective potential, we take $S_{3}$ to be
@@
We can now formulate the system of coupled first order ODEs which describe the background evolution in the false vacuum during the FOPT. The ODEs track $x\equiv H_{\rm c}(t-t_{\rm c}),y\equiv T/T_{\rm c},y_{\rm SM}\equiv T_{\rm SM}/T_{\rm SM,c},F$, the derivatives of $F$, and $\tilde{\eta}\equiv\mathcal{H}_{\rm c}(\eta-\eta_{\rm c})$. We set $a=1$ at the critical point. This then implies that $H_{\rm c}=\mathcal{H}_{\rm c}$. Taking the scale factor $a$ to be the independent variable, one can show that
@@
$g_{n}(t)\equiv-\frac{4\pi}{3}~\frac{3!}{(3-n)!}H_{\rm c}^{4-n}\int_{t_{\rm c}}^{t}dt^{\prime}~\frac{\Gamma(T(t^{\prime}))}{H_{\rm c}^{4}}a^{3}(t^{\prime})~r^{3-n}(t,t^{\prime})\,.$ (4.35)
The $g_{n}$s are dimensionless quantities that are related to the derivatives of $\ln F$, and can also be defined recursively via Eq. (4.32). A similar system of ODEs, in particular that converts the integro-differential equation into a system of ODEs, was obtained in Ref. [kampfer1988phenomenologicalPaper1] and recently in Ref. [Flores:2024lng]. Note that our system of ODEs has one less equation than Ref. [Flores:2024lng], if we do not include Eq. (4.34) in the counting, which only tracks the elapsed conformal time from the critical point in units of $H_{\rm c}$. The initial conditions are
@@
for $a=1$. We check that the physical volume in the false vacuum regions is decreasing, at least at the percolation time. This can be expressed as [Turner:1992tz, Athron:2022mmm]
@@
Figure 3: Evolution of the dark/visible sector temperatures and FV fraction (top row), and of $c_{s\rm,D}^{2}$ and $w_{\rm D}$ (bottom row), for BP-2 (column a) and BP-4 (column b). The vertical black dashed lines indicate the percolation temperature in all panels, while the vertical dot-dashed and dotted lines respectively indicate $\tilde{\eta}_{i}$ and $\tilde{\eta}_{f}$ in the bottom panels.
@@
For illustration purposes, in Fig. 3 and in the following figures, we feature two specific FOPT scenarios: BP-2 and BP-4 from Table 1. For each case, the evolution of the temperature (solid) and false vacuum fraction (dashed) versus the rescaled elapsed conformal time $H_{\rm c}(\eta-\eta_{\rm c})$, are shown in the top row plots. The labeling is based on Table 1, where we list the benchmark FOPT scenarios of interest.
BP-1 BP-2 BP-3 BP-4 BP-5 BP-6
AA $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-4}$ $10^{-4}$ $10^{-4}$
$x_{C}$ $1.1063\times 10^{-4}$ $2.4432\times 10^{-4}$ $2.4131\times 10^{-4}$ $1.712\times 10^{-5}$ $3.97\times 10^{-7}$ $8.19\times 10^{-6}$
$\lambda$ $3.6917\times 10^{-4}$ $3.6917\times 10^{-4}$ $4.8063\times 10^{-4}$ $1.7136\times 10^{-5}$ $1.7136\times 10^{-5}$ $2.2309\times 10^{-5}$
$L_{\rm c}/T_{\rm c}^{4}$ $5.0456\times 10^{-2}$ $9.0072\times 10^{-2}$ $3.998\times 10^{-2}$ $7.2479\times 10^{-2}$ $2.8264\times 10^{-2}$ $1.9733\times 10^{-2}$
$T_{\rm c}$ (MeV) $10^{3}$ $^{-2} 1010$ $^{3} 1010$ $10^{5}$
$r_{T\rm,c}$ 0.4 0.4 0.4 0.4 0.1 0.4
$T_{*}$ (MeV) $3.4306\times 10^{2}$ $2.3654\times 10^{-3}$ 2.4216 $3.0246\times 10^{2}$ 3.8815 $3.3936\times 10^{4}$
$T_{\rm SM,*}$ (MeV) $8.9392\times 10^{2}$ $5.9135\times 10^{-3}$ 6.2448 $7.9351\times 10^{2}$ $4.1986\times 10^{1}$ $8.5999\times 10^{4}$
$\Delta N_{\rm eff}$ $3.4901\times 10^{-3}$ 0.2496 $4.7392\times 10^{-2}$ $3.4165\times 10^{-3}$ $1.0787\times 10^{-4}$ $2.5327\times 10^{-3}$
$\sqrt{3}H_{*}R_{*}$ $1.1716\times 10^{-3}$ $1.7169\times 10^{-3}$ $2.3446\times 10^{-3}$ $9002\times 10^{-4} 2.2$$2558\times 10^{-4}$ $3.4178\times 10^{-4}$
$M_{*}/M_{\odot}$ $2.88\times 10^{-14}$ 0.2359 $8.73\times 10^{-8}$ $5.44\times 10^{-16}$ $3.8\times 10^{-15}$ $5.3\times 10^{-20}$
$\alpha_{*}$ $1.2\times 10^{-5}$ $8.1792\times 10^{-4}$ $3.19\times 10^{-5}$ $5.1\times 10^{-5}$ $2.71\times 10^{-7}$ $4.99\times 10^{-6}$
$\beta_{*}/H_{*}$ $2.2962\times 10^{3}$ $5653\times 10^{3} 1.1$$1432\times 10^{3}$ $9.309\times 10^{3}$ $1.3535\times 10^{4}$ $8.1307\times 10^{3}$
$v_{\rm w,*}$ 0.9998 1 1 0.9999 0.9995 0.9998
$s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$ 7.3349 0.1379 0.2581 $2304\times 10^{2} 1.1$$1132\times 10^{4}$ $1.0651\times 10^{2}$
$\tilde{\eta}_{*}-\tilde{\eta}_{i}$ $1.2187\times 10^{-2}$ $2.5631\times 10^{-2}$ $3.2774\times 10^{-2}$ $3.8714\times 10^{-3}$ $2.08\times 10^{-3}$ $3.9741\times 10^{-3}$
$\tilde{\eta}_{*}$ 1.8899 3.2276 3.1085 2.2708 1.5419 1.9378
$c_{s\rm,D*}^{2}$ -$1.664\times 10^{1}$ -$4.0616\times 10^{1}$ -$4.0832\times 10^{1}$ -$1.3081\times 10^{1}$ -6.9385 -$1.4892\times 10^{1}$
$\mathcal{C}_{*}$ 0.8866 $3.555\times 10^{-2}$ $1.866\times 10^{-3}$ $3.0621\times 10^{1}$ $5.5423\times 10^{1}$ 2.2367
$v_{\rm eff,*}/A_{\rm s}$ $8.2735\times 10^{-4}$ $2.4278\times 10^{-4}$ $7.6\times 10^{-5}$ $1.2037\times 10^{-3}$ $1.2595\times 10^{-3}$ $3.8337\times 10^{-4}$
$k_{\rm cut}/\mathcal{H}_{\rm c}$ $5.6035\times 10^{2}$ $1.7729\times 10^{2}$ $1.3711\times 10^{2}$ $1.7669\times 10^{3}$ $3.7758\times 10^{3}$ $1.8512\times 10^{3}$
$k_{\rm th}/\mathcal{H}_{\rm c}$ $7.3322\times 10^{2}$ $1.7558\times 10^{2}$ $1.8489\times 10^{2}$ $1.7493\times 10^{3}$ $5.9323\times 10^{3}$ $2.5752\times 10^{3}$
$k_{\rm cut}/k_{\rm th}$ 1.3085 0.9904 1.3485 0.9900 1.5711 1.3911
$s_{\rm rms}$ $6.0685\times 10^{-3}$ $3.3471\times 10^{-5}$ $1.96\times 10^{-5}$ 0.1481 $1.4021\times 10^{1}$ $4.0831\times 10^{-2}$
@@
We mark the moment of percolation with a vertical dashed line. In both cases, we have taken $r_{T\rm,c}=0.4$. For these two cases, the percolation temperatures are such that $T_{*}/T_{\rm c}<0.9$, and the phase transition strength $\alpha_{*}$ is at least $10^{-7}$. Here we define $\alpha_{*}$ as
$\alpha_{*}$ $\equiv$ $\frac{1}{\rho_{\rm R,tot}}\left(1-T\frac{d}{dT}\right)\Delta V(T)\Bigg|_{T=T_{*}},\quad\rho_{\rm R,tot}(T_{*})=\frac{\pi^{2}}{30}T_{*}^{4}\left[g_{\rho}+\frac{g_{\rho,{\rm SM}}(T_{*})}{r_{T\rm,*}^{4}}\right]\,.$ (4.39)
The conformal time elapsed from the critical point to the percolation time can be estimated by first noting that $T\propto 1/a$ and $a\propto\eta$ in the radiation-dominated era. Then
@@
From the above estimate, a lower percolation temperature implies a longer conformal time duration between the critical point and percolation. The percolation temperature, in turn, can be estimated from Eq. (4.22) by approximating
$\Gamma(T(t^{\prime}))\approx\Gamma(t)\exp\left[\beta(t)(t-t^{\prime})\right],\quad\frac{\beta(t)}{H}\equiv\frac{1}{H}\frac{\dot{\Gamma}}{\Gamma}\simeq T\frac{d}{dT}\left(\frac{S_{3}}{T}\right)\,$, (4.41)
where $\beta(t)$ is the instantaneous inverse time scale of the FOPT at time $t$. Assuming that the scale factor is approximately constant and $\beta(t)$ varies slowly within the period of interest, we have [Enqvist:1991xw]
$-\ln F(t)$ $\approx 8\pi\frac{v_{\rm w}^{3}}{\beta^{4}(t)}\Gamma(t)\,.$ (4.42)
At percolation, we then have
$1\approx 8\pi v_{\rm w,*}^{3}\frac{\Gamma_{*}}{\beta_{*}^{4}}\Rightarrow\exp\left[\frac{S_{3}(T_{*})}{T_{*}}\right]\approx 8\pi\left[\frac{S_{3}(T_{*})}{T_{*}}\right]^{3/2}\frac{v_{\rm w,*}^{3}}{(\beta_{*}/H_{*})^{4}}\left(\frac{T_{*}}{H_{*}}\right)^{4}\,.$ (4.43)
Typically, the bounce action at percolation for an FOPT with $T_{\rm c}=10\,\mathrm{MeV}$ and $r_{T\rm,c}=0.4$ is $S_{3}/T\sim 150$. In Fig. 4 we plot $S_{3}/T$ as solid curves, where we feature BP-2 (red) and BP-4 (blue). These benchmark points correspond to critical temperatures of 10 keV and 1 GeV, respectively. The actual values of $S_{3}(T_{*})/T_{*}$ for BP-2 and BP-4 are, respectively, 194.7 and 133.5. In general, the bounce action can change depending on the scale of $T_{\rm c}$. Based on the estimate in Eq. (4.43), the bounce action can increase, relative to the value at $T_{\rm c}=10\,\mathrm{MeV}$, by $4\ln\left[\left(T_{\rm c}/10\,\mathrm{MeV}\right)\right]$, where we approximated $T_{*}\simeq T_{\rm c}$.
Figure 4: $S_{3}/T$ and $\beta/H$, for BP-2 (red) and BP-4 (blue). The vertical dot-dashed lines indicate the percolation temperature for each benchmark point.
In order to account for the presence of the FV and TV, the thermodynamic properties of the dark fluid are determined by the average of the FV and TV contributions, weighted by the respective volume fractions, i.e.,
@@
The above prescription, which uses volume fractions as weights to account for the contribution of each phase, is valid over sufficiently large scales where FV and TV features are indistinguishable. The dark sector equation of state and the sound speed are then respectively given by
$\boxed{w_{\rm D}=\frac{p_{\rm D}}{\rho_{\rm D}},\quad c_{s\rm,D}^{2}=\frac{dp_{\rm D}/dt}{d\rho_{\rm D}/dt}=\frac{s_{\rm D}-\left(\frac{\dot{F}}{H}\right)\frac{\Delta p}{T}}{T~\frac{\partial s_{\rm D}}{\partial T}-\left(\frac{\dot{F}}{H}\right)\frac{\Delta\rho}{T}}}\,.$ (4.45)
@@
During the short period when the false vacuum fraction drops rapidly, i.e., when the system quickly transforms from the false to the true vacuum configuration, we have $|\dot{F}|/H\gg 1$, and
$c_{s\rm,D}^{2}\approx\frac{\Delta p}{\Delta\rho}.$ (4.47)
Note that the sound speed squared can be negative since $\Delta p$ and $\Delta\rho$ have opposite signs when $T<T_{\rm c}$. Sufficiently short wavelength perturbations whose scales are comparable with the duration of the sudden drop in $F$, are expected to grow. We emphasize that $c_{s,i}^{2}$ in this limit refers to the sound speed in the FV or the TV, when each is taken in isolation. This is the sound speed that has been computed in, e.g., Refs. [Espinosa:2010hh, Tenkanen:2022tly]. However, during the phase transition, the sound speed $c_{s\rm,D}^{2}$ in the dark plasma may deviate from the value suggested by $c_{s\rm,FV}^{2}=1/3$, or $c_{s\rm,TV}^{2}$, not necessarily equal to 1/3. For example, if the false vacuum fraction decreases sharply, the denominator in Eq. (4.45) can be large so that $c_{s\rm,D}^{2}$ can become negative. A similar situation occurs in the physical system considered in Ref. [Schmid:1998mx], where the system undergoes prolonged phase coexistence at $T\simeq T_{\rm c}$, keeping the pressure constant, which leads to a zero sound speed in the long-wavelength limit. On the other hand, Ref. [Tenkanen:2022tly] mentioned in a footnote that the sound speed does not vanish at the critical point, which is contrary to the result in Ref. [Schmid:1998mx]. This is not very surprising, since the former considered the sound speeds in each phase taken in isolation, i.e., $c_{s,\rm FV}^{2}$ and $c_{s\rm,TV}^{2}$. For the sake of clarity, in Table 2 we explicitly define quantities, such as the equation of state and sound speed squared, that refer to the dark sector; SM sector; and the average over the dark and SM sectors.
@@
$p_{\rm D},\rho_{\rm D}$ Dark sector pressure, density averaged over the false and true vacua
$w_{\rm D}$ Dark sector equation of state
@@
$w$ Equation of state, averaged over the dark and visible sectors
$c_{s\rm,D}^{2}$ Dark sector sound speed squared